When a clock hand measuring cm in length moves through 40 minutes, how far does the tip move? Give your answer correct to 2 decimal places.
Answer by checkley77(12569)

the circumferance of a circle with a radius of cm =2*pi*r=2*3.14*cm==6.28*cm
40/60*6.28*cm=2/3*6.28*cm=4.19*cm answer.